Vulcan, Inc., has 8 percent coupon bonds on the market that have 8 years left to maturity. The bonds make annual payments and have a par value of $1,000.

If the YTM on these bonds is 10 percent, what is the current bond price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Annual coupon=1000*8%=80

Hence current price=Annual coupon*Present value of annuity factor(10%,8)+$1000*Present value of discounting factor(10%,8)

=80*5.3349262+$1000*0.46650738

=$893.30(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=80[1-(1.1)^-8]/0.1

=80*5.3349262

2.Present value of discounting factor=1000/1.1^8

=$1000*0.46650738